# Geometric Stability: The Missing Axis of Representations

### v1

@@ 3.1 Stability Predicts Steering Performance
To isolate the contribution of stability from mere class separation, we computed partial correlations to control for the Fisher discriminant and silhouette score. Shesha maintained large effect sizes across all settings: $\rho_{\text{partial}}=0.67$ (Synthetic), $0.76$ (SST-2), and $0.62$ (MNLI), all $p<0.001$. This demonstrates that the reliability of class structure under perturbation, geometric consistency, is a distinct factor of steerability that separability metrics overlook. Separability may be necessary for steering, but stability ensures control.
@@ 3.2 Visual Perception: The Architecture of Stability
We evaluated 94 pretrained vision models that encompassed diverse architectures (ViT, Swin, ConvNeXt, ResNet, EfficientNet) and training objectives (supervised, self-supervised, contrastive, generative) across six datasets in four visual domains: CIFAR-10/100 (natural images), Flowers-102 and Oxford Pets (fine-grained), DTD (texture), and EuroSAT (remote sensing). All metrics were computed on the features of the penultimate-layer (Appendix 9).
@@
On CIFAR-10, Shesha showed a negligible correlation with LogME ($\rho=-0.002$). However, as the task complexity increased, the relationship inverted: CIFAR-100 showed a negative trend ($\rho=-0.17$), while Oxford Pets showed the strongest negative correlation ($\rho=-0.44$, $p<0.001$). The stability-transferability tradeoff intensifies with both class count and demands for fine-grained discrimination. EuroSAT was the sole exception, showing a positive correlation ($\rho=0.45$, $p<0.001$), possibly due to uniform texture distributions.
The DINO Paradox: universal across domains.
The stability-transferability inversion persisted across all six datasets (Figure 15). DINOv2 achieved the highest mean LogME on 4/6 datasets but ranked last or near-last in geometric stability on all except EuroSAT. On CIFAR-100: LogME =1.36=1.36 (rank 1/19 families), Shesha =0.27=0.27 (rank 18/19). On Flowers-102: LogME =2.47=2.47 (rank 1/19), Shesha =0.34=0.34 (rank 19/19). On CIFAR-10, DINOv2 ranked first in LogME but last in stability (rank 19/19). Optimizing for state-of-the-art transfer appears to incur a “geometric tax” by collapsing the manifold structure required for stability. The sole exception was EuroSAT, where DINOv2 achieved top stability (Shesha =0.95=0.95, rank 3/19) while also ranking first in LogME. This suggests that DINOv2’s self-distillation objective may be uniquely suited to satellite imagery geometry.
@@
Two factors consistently predicted geometric stability: contrastive alignment and hierarchical architecture. CLIP models significantly outperformed self-supervised alternatives in 4 out of 6 domains ($p<0.01$), and EVA-02 achieved the highest stability by reconstructing CLIP features instead of pixels. This confirms that alignment targets determine geometry. Swin, PVT, and CoAtNet significantly outperformed columnar architectures (ViT, DeiT) on 5 out of 6 datasets, with multi-scale processing acting as implicit geometric regularization.
@@
CIFAR-10/100 rankings correlated at $\rho=0.92$, confirming that stability is intrinsic rather than task-specific. However, Flowers-102 vs. Oxford Pets showed no correlation ($\rho=-0.03$), suggesting fine-grained domains require distinct stability properties. Consistent patterns emerged across datasets: CLIP, EVA, and Swin families maintained high stability, while DINOv2, DeiT3, and ViT families showed low stability (except DINOv2 on EuroSAT).
@@ 4 Discussion
Our vision architecture analysis reveals a fundamental tension between transfer learning and geometric stability, modulated by task complexity. On simple tasks (CIFAR-10) or uniform imagery (EuroSAT), most models perform adequately on both axes. However, as class count increases or fine-grained discrimination is required, specialization becomes necessary. DINOv2 exemplifies this trade-off: it achieves state-of-the-art transferability (LogME rank 1/19 on 4/6 datasets) while ranking last or near-last in geometric stability on all but EuroSAT. This “geometric tax” suggests that optimizing for rich, adaptable features collapses the manifold structure required for predictable behavior. For practitioners, this implies distinct model selection criteria: high-LogME models (DINOv2, ResNets) for transfer learning pipelines; high-Shesha models (CLIP, EVA-02, Swin) for zero-shot or safety-critical deployment where behavioral consistency matters more than adaptability.
@@
Two factors consistently predict geometric stability across visual domains. First, contrastive alignment: CLIP models significantly outperformed self-supervised alternatives on 4/6 datasets, and EVA-02 achieved maximal stability by reconstructing CLIP features rather than pixels, confirming that alignment targets, not training mechanisms, determine geometry. Second, hierarchical architecture: Swin, PVT, and CoAtNet demonstrated superior performance to columnar architectures (ViT, DeiT) on 5/6 datasets, which indicates that multi-scale processing acts as implicit geometric regularization. These findings can assist practitioners in selecting architectures. When stability is required, prefer contrastively-aligned models with hierarchical feature pyramids.
@@
Our results across steering, transfer learning, and vision reveal a fundamental boundary condition for geometric stability. In domains where the task is intrinsic to the data manifold, such as drift detection and biological fitness, unsupervised stability is highly predictive. However, in high-level semantic tasks, intrinsic stability is insufficient. We observed an important dissociation in Real-World Steering (Section 3.1): supervised Shesha achieved near-perfect prediction ($\rho\geq 0.96$), proving that stability drives control, whereas unsupervised Shesha failed entirely ($\rho\leq 0.35$, n.s.); in Vision (Section 3.2), stability and transferability were empirically independent ($\rho\approx 0$) or negatively correlated on complex tasks ($\rho=-0.17$ to 0.44-0.44). This dissociates “rigidity” from “utility”: a model can be geometrically rigid yet semantically misaligned (DINOv2: top transfer, bottom stability), necessitating either supervised stability metrics or explicit consideration of task alignment when predicting downstream control.
@@ 5 Related Work
RLHF and related alignment methods are known to induce an “alignment tax” on broader capabilities (Ouyang et al.,, 2022; Bai et al.,, 2022; Lin et al.,, 2024). Studies have identified representation collapse under preference optimization (Moalla et al.,, 2024) and mode collapse in RLHF (O’Mahony et al.,, 2024) through behavioral metrics. Geometry remains less explored, though emerging evidence shows distributional and feature-space drift (Li et al., 2025b, ). Our analysis provides the first systematic quantification of post-alignment geometric drift (Section LABEL:sec:app-drift).
@@ 9 Visual Perception Architecture: Extended Methods and Results
This appendix provides complete experimental details and supplementary analyses for the visual architecture benchmark described in Section 3.2. We evaluate geometric stability as a diagnostic for transfer learning through systematic comparison of 94 pretrained vision models across six datasets. The first analysis established the relationship between stability (Shesha-FS) and transferability (LogME, LEEP) across natural image domains (CIFAR-10, CIFAR-100). The second tested generalization to fine-grained recognition (Flowers-102, Oxford Pets), texture classification (DTD), and remote sensing (EuroSAT). The third characterized architecture family differences in geometric stability profiles.
@@ 9.2.1 Correlation Structure Across Datasets
LogME vs Shesha-Var $0.67^{***}$ $0.50^{***}$ $0.30^{**}$ $4^{*} 0.220$$2^{*}$ $0.30^{**}$
LogME vs Shesha-FS 0.00-0.00 0.17-0.17 0.08-0.08 $0.19$ $45^{***} 0.0$70.07
Shesha-Var vs Shesha-FS 0.12-0.12 $25^{*} 0.0$90.09 $0.03$ 0.11-0.11 $0.44^{***}$
@@ 9.2.3 The DINOv2 Paradox
Table 25: The DINOv2 Paradox Across Datasets. DINOv2 achieves highest LogME but lowest Shesha-FS on every dataset except EuroSAT.
@@
CIFAR-10 1.020 1/27 0.378 27/27
CIFAR-100 1.360 1/27 0.273 27/27
Flowers-102 2.466 1/27 0.337 27/27
DTD 0.878 4/27 0.466 22/27
EuroSAT 0.572 3/27 0.948 1/27
Oxford Pets 1.280 4/27 0.530 24/27
@@
Table 26: Contrastive vs. Self-Supervised Stability (Shesha-FS). Mann-Whitney U tests comparing CLIP ($n=3$) to SSL models ($n=10$).
@@
CIFAR-10 $0.84\pm 0.04$ $0.55\pm 0.16$ +0.29 $0.003^{**}$
CIFAR-100 $0.83\pm 0.06$ $0.47\pm 0.18$ +0.35 $0.003^{**}$
Flowers-102 $0.86\pm 0.02$ $0.64\pm 0.33$ +0.22 $0.346$
DTD $0.76\pm 0.03$ $0.50\pm 0.07$ +0.26 $0.003^{**}$
EuroSAT $0.95\pm 0.01$ $0.86\pm 0.10$ +0.10 $0.143$
Oxford Pets $0.85\pm 0.01$ $0.55\pm 0.13$ +0.30 $0.007^{**}$
$^{**}p<0.01$
@@
Dataset Hierarchical ($n=16$) Columnar ($n=13$) $\Delta$ $p$
CIFAR-10 $0.71\pm 0.07$ $0.53\pm 0.12$ +0.18 $0.001^{***}$
CIFAR-100 $0.64\pm 0.07$ $0.41\pm 0.15$ +0.23 $0.001^{***}$
Flowers-102 $0.92\pm 0.04$ $0.74\pm 0.20$ +0.19 $0.001^{***}$
DTD $0.51\pm 0.06$ $44\pm 0.09 +0.0$7 $0.017^{*}$
EuroSAT $0.83\pm 0.05$ $74\pm 0.18 +0.0$9 $0.046^{*}$
Oxford Pets $0.59\pm 0.08$ $57\pm 0.07 +0.0$2 $0.439$
$^{*}p<0.05$; $^{***}p<0.001$
@@ 9.2.7 Architecture Family Stability Profiles
Table 29: Complete Family Shesha-FS Scores Across Datasets. Ordered by mean cross-dataset stability.
@@ 9.2.8 Top-20 Rank Agreement
LEEP vs LogME 0.54 0.54 0.14 0.29 0.25 0.29
LEEP vs Shesha-Var 0.60 0.48 0.29 0.11 0.21 0.11
LEEP vs Shesha-FS 0.14 0.11 0.14 0.11 0.11 0.14
LogME vs Shesha-Var 0.54 0.54 0.21 0.03 0.18 0.11
LogME vs Shesha-FS 0.14 0.11 0.08 0.25 0.18 0.25
Shesha-Var vs Shesha-FS 0.11 0.05 0.25 0.11 0.14 0.14
Models achieving top-20 status for all four metrics: CIFAR-10 (3 models: Swin-Large, ConvNeXt-Base, ConvNeXtV2-Base), CIFAR-100 (1 model: Swin-Large), Flowers-102 (1 model: ResNeXt-101), DTD (1 model: ResNeXt-101), EuroSAT (1 model: RegNetY-032), Oxford Pets (1 model: ResNeXt-101). The weakly-supervised ResNeXt-101 appears as the only “universal generalist” across 4/6 datasets.
@@ 9.2.9 Statistical Tests Summary
CIFAR-10 $34.4^{***}$ $p<0.001^{***}$ $p=0.016^{*}$ $p=0.003^{**}$
CIFAR-100 $40.7^{***}$ $p<0.001^{***}$ $p=0.040^{*}$ $p=0.003^{**}$
Flowers-102 $35.5^{***}$ $^{***} p<0.001p<0.001^{***}$ $p=0.346$
DTD $34.0^{***}$ $p=0.017^{*}$ $p=0.635$ $p=0.003^{**}$
EuroSAT $23.5^{**}$ $p=0.046^{*}$ $p=0.028^{*}$ $p=0.143$
Oxford Pets $38.2^{***}$ $p=0.439$ $p=0.226$ $p=0.007^{**}$

### v2

@@ 3.1 Stability Predicts Steering Performance
The raw correlations were consistently strong: $\rho=0.89$ (Synthetic), $0.96$ (SST-2), and $0.96$ (MNLI), all $p<10^{-18}$. To isolate the contribution of stability from mere class separation, we computed partial correlations to control for the Fisher discriminant and silhouette score. Shesha maintained large effect sizes across all settings: $\rho_{\text{partial}}=0.67$ (Synthetic), $0.76$ (SST-2), and $0.62$ (MNLI), all $p<0.001$. This demonstrates that the reliability of class structure under perturbation, geometric consistency, is a distinct factor of steerability that separability metrics overlook. Separability may be necessary for steering, but stability ensures control.
@@ 3.2 Visual Perception: The Architecture of Stability
We evaluated 93-94 pretrained vision models that encompassed diverse architectures (ViT, Swin, ConvNeXt, ResNet, EfficientNet) and training objectives (supervised, self-supervised, contrastive, generative) across six datasets in four visual domains: CIFAR-10/100 (natural images), Flowers-102 and Oxford Pets (fine-grained), DTD (texture), and EuroSAT (remote sensing). All metrics were computed on the features of the penultimate-layer (Appendix 9)
@@
On CIFAR-10, Shesha-FS showed a negligible correlation with LogME ($\rho=-0.07$). However, as the task complexity increased, the relationship varied by domain: CIFAR-100 showed a negative trend ($\rho=-0.19$), and Flowers-102 showed a significant negative correlation ($\rho=-0.21$, $p<0.05$). Oxford Pets exhibited the strongest negative correlation between Shesha-Var and Shesha-FS ($\rho=-0.42$, $p<0.001$), indicating that models with higher variance showed lower stability on fine-grained tasks. EuroSAT was a notable exception, showing a positive LogME-Shesha-FS correlation ($\rho=0.45$, $p<0.001$), possibly due to uniform texture distributions in satellite imagery.
The DINOv2 Paradox: universal across domains.
The stability-transferability inversion persisted across all six datasets (Figure 15). DINOv2 achieved the highest mean LogME on 4/6 datasets but ranked last or near-last in geometric stability on all except EuroSAT. On CIFAR-100: LogME =1.36=1.36 (rank 1/29 families), Shesha-FS =0.27=0.27 (rank 28/29). On Flowers-102: LogME =2.47=2.47 (rank 1/29), Shesha-FS =0.34=0.34 (rank 29/29). On CIFAR-10, DINOv2 ranked first in LogME but last in stability (rank 29/29). Optimizing for state-of-the-art transfer appears to incur a “geometric tax” by collapsing the manifold structure required for stability. The sole exception was EuroSAT, where DINOv2 achieved top stability (Shesha-FS =0.95=0.95, rank 3/29) while also ranking second in LogME (rank 2/29). This suggests that DINOv2’s self-distillation objective may be uniquely suited to satellite imagery geometry.
@@
Two factors consistently predicted geometric stability: contrastive alignment and hierarchical architecture. CLIP models outperformed self-supervised alternatives in 4 out of 6 domains ($p<0.05$), and EVA-02 achieved the highest stability by reconstructing CLIP features instead of pixels. This confirms that alignment targets determine geometry. Hierarchical transformers (Swin, PVT, MaxViT, CoAtNet, PoolFormer) significantly outperformed columnar architectures (ViT, DeiT) on CIFAR-10 ($p=0.011$), CIFAR-100 ($p=0.007$), and Flowers-102 ($p<0.001$), with multi-scale processing acting as implicit geometric regularization. However, this advantage did not generalize to DTD, EuroSAT, or Oxford Pets.
@@
CIFAR-10/100 rankings correlated at $\rho=0.92$, confirming that stability is intrinsic rather than task-specific. However, Flowers-102 vs. Oxford Pets showed no correlation ($\rho=-0.03$), suggesting fine-grained domains require distinct stability properties. Consistent patterns emerged across datasets: CLIP, EVA, and Inception families maintained high stability, while DINOv2, DeiT3, and ViT families showed low stability (except DINOv2 on EuroSAT).
@@ 4 Discussion
Our vision architecture analysis reveals a fundamental tension between transfer learning and geometric stability, modulated by task complexity. On simple tasks (CIFAR-10) or uniform imagery (EuroSAT), most models perform adequately on both axes. However, as class count increases or fine-grained discrimination is required, specialization becomes necessary. DINOv2 exemplifies this trade-off: it achieves state-of-the-art transferability (LogME rank 1/29 on 4/6 datasets) while ranking last or near-last in geometric stability on all but EuroSAT. This “geometric tax” suggests that optimizing for rich, adaptable features collapses the manifold structure required for predictable behavior. For practitioners, this implies distinct model selection criteria: high-LogME models (DINOv2, ResNets) for transfer learning pipelines; high-Shesha-FS models (CLIP, EVA-02, Inception) for zero-shot or safety-critical deployment where behavioral consistency matters more than adaptability.
@@
Two factors consistently predict geometric stability across visual domains. First, contrastive alignment: CLIP models outperformed self-supervised alternatives on 4/6 datasets ($p<0.05$), and EVA-02 achieved maximal stability by reconstructing CLIP features rather than pixels, confirming that alignment targets, not training mechanisms, determine geometry. Second, hierarchical architecture: Swin, PVT, and CoAtNet demonstrated superior performance to columnar architectures (ViT, DeiT) on 3/6 datasets (CIFAR-10, CIFAR-100, Flowers-102), which indicates that multi-scale processing acts as implicit geometric regularization, though this advantage did not generalize to texture, satellite, or fine-grained animal domains. These findings can assist practitioners in selecting architectures. When stability is required, prefer contrastively-aligned models with hierarchical feature pyramids.
@@
Our results across steering, transfer learning, and vision reveal a fundamental boundary condition for geometric stability. In domains where the task is intrinsic to the data manifold, such as drift detection and biological fitness, unsupervised stability is highly predictive. However, in high-level semantic tasks, intrinsic stability is insufficient. We observed an important dissociation in Real-World Steering (Section 3.1): supervised Shesha achieved near-perfect prediction ($\rho\geq 0.96$), proving that stability drives control, whereas unsupervised Shesha failed entirely ($\rho\leq 0.35$, n.s.); in Vision (Section 3.2), stability and transferability were empirically independent ($\rho\approx 0$) or negatively correlated on complex tasks ($\rho=-0.19$ to 0.42-0.42). This dissociates “rigidity” from “utility”: a model can be geometrically rigid yet semantically misaligned (DINOv2: top transfer, bottom stability), necessitating either supervised stability metrics or explicit consideration of task alignment when predicting downstream control.
@@ 5 Related Work
RLHF and related alignment methods are known to induce an “alignment tax” on broader capabilities (Ouyang et al.,, 2022; Bai et al.,, 2022; Lin et al.,, 2024). Studies have identified representation collapse under preference optimization (Moalla et al.,, 2024) and mode collapse in RLHF (O’Mahony et al.,, 2024) through behavioral metrics. Geometry remains less explored, though emerging evidence shows distributional and feature-space drift (Li et al., 2025b, ). Our analysis provides the first systematic quantification of post-alignment geometric drift (Section 3.3).
@@ 9 Visual Perception Architecture: Extended Methods and Results
This appendix provides complete experimental details and supplementary analyses for the visual architecture benchmark described in Section 3.2. We evaluate geometric stability as a diagnostic for transfer learning through systematic comparison of 93-94 pretrained vision models across six datasets. The first analysis established the relationship between stability (Shesha-FS) and transferability (LogME, LEEP) across natural image domains (CIFAR-10, CIFAR-100). The second tested generalization to fine-grained recognition (Flowers-102, Oxford Pets), texture classification (DTD), and remote sensing (EuroSAT). The third characterized architecture family differences in geometric stability profiles.
@@ 9.2.1 Correlation Structure Across Datasets
LogME vs Shesha-Var $9^{***} 0.550$$5^{***}$ $47^{***} 0.0$3-0.03 $0.33^{**}$ $0.36^{***}$
LogME vs Shesha-FS 0.07-0.07 0.19-0.19 $0.21^{*}$ $0.17$ $45^{***} 0.0$8-0.08
Shesha-Var vs Shesha-FS 0.01-0.01 $0.24^{*}$ 0.14-0.14 $0.06$ $0.12$ $0.42^{***}$
@@ 9.2.3 The DINOv2 Paradox
Table 25: The DINOv2 Paradox Across Datasets. DINOv2-giant achieves highest LogME on 4/6 datasets while consistently ranking in the bottom quartile for Shesha-FS, except on EuroSAT where it ranks first on both metrics.
@@
CIFAR-10 1.386 1/94 0.415 88/94
CIFAR-100 1.629 1/94 0.319 86/94
Flowers-102 3.521 1/93 0.152 92/93
DTD 0.952 19/93 0.502 62/93
EuroSAT 0.681 1/93 0.987 1/93
Oxford Pets 1.760 19/93 0.569 68/93
Table 26: The DINOv2 Paradox (Family Averages). DINOv2 family achieves highest LogME on multiple datasets while ranking last or near-last in Shesha-FS, except on EuroSAT.
Dataset LogME LogME Rank Shesha-FS FS Rank
CIFAR-10 1.020 1/29 0.378 29/29
CIFAR-100 1.360 1/29 0.273 28/29
Flowers-102 2.466 1/29 0.337 29/29
DTD 0.878 9/29 0.466 24/29
EuroSAT 0.572 2/29 0.948 3/29
Oxford Pets 1.280 9/29 0.530 25/29
@@
Table 27: Contrastive vs. Self-Supervised Stability (Shesha-FS). Mann-Whitney U tests comparing CLIP ($n=3$) to SSL models ($n=9$).
@@
CIFAR-10 $0.84\pm 0.04$ $0.57\pm 0.20$ +0.27 $0.032^{*}$
CIFAR-100 $0.83\pm 0.06$ $0.48\pm 0.24$ +0.34 $0.032^{*}$
Flowers-102 $0.86\pm 0.02$ $0.54\pm 0.28$ +0.32 $0.032^{*}$
DTD $0.76\pm 0.03$ $0.55\pm 0.13$ +0.21 $0.032^{*}$
EuroSAT $0.95\pm 0.01$ $91\pm 0.08 +0.0$4 $0.568$
Oxford Pets $0.85\pm 0.01$ $0.68\pm 0.19$ +0.17 $0.141$
$^{*}p<0.05$; $^{**}p<0.01$; $^{***}p<0.001$
@@
Dataset Hierarchical ($n=18$) Columnar ($n=23$) $\Delta$ $p$
CIFAR-10 $0.70\pm 0.07$ $0.58\pm 0.18$ +0.12 $0.011^{*}$
CIFAR-100 $0.63\pm 0.07$ $0.48\pm 0.23$ +0.15 $0.007^{**}$
Flowers-102 $0.92\pm 0.04$ $0.66\pm 0.24$ +0.26 $0.001^{***}$
DTD $0.51\pm 0.06$ $52\pm 0.15 -0.0$1 $0.197$
EuroSAT $0.83\pm 0.04$ $83\pm 0.17 +0.0$0 $0.844$
Oxford Pets $0.59\pm 0.08$ $65\pm 0.16 -0.0$7 $0.941$
$^{*}p<0.05$; $^{**}p<0.01$; $^{***}p<0.001$
@@ 9.2.7 Architecture Family Stability Profiles
Table 30: Family Shesha-FS Scores Across Datasets. Ordered by mean cross-dataset stability.
@@ 9.2.8 Top-20 Rank Agreement
LEEP vs LogME 0.43 0.43 0.11 0.29 0.14 0.29
LEEP vs Shesha-Var 0.38 0.29 0.18 0.05 0.11 0.11
LEEP vs Shesha-FS 0.08 0.05 0.14 0.11 0.05 0.11
LogME vs Shesha-Var 0.33 0.29 0.60 0.05 0.21 0.11
LogME vs Shesha-FS 0.11 0.05 0.05 0.18 0.29 0.14
Shesha-Var vs Shesha-FS 0.14 0.11 0.05 0.14 0.29 0.08
Models achieving top-20 status for all four metrics: CIFAR-10 (1 model: Swin-Large), CIFAR-100 (0 models), Flowers-102 (1 model: ResNeXt-101), DTD (1 model: ResNeXt-101), EuroSAT (1 model: RegNetY-032), Oxford Pets (0 models). The weakly-supervised ResNeXt-101 appears as a “universal generalist” on 2/6 datasets (Flowers-102, DTD).
@@ 9.2.9 Statistical Tests Summary
CIFAR-10 $58.3^{***}$ $p=0.001^{**}$ $p=0.867$ $p=0.013^{*}$
CIFAR-100 $63.3^{***}$ $p<0.001^{***}$ $p=0.996$ $p=0.013^{*}$
Flowers-102 $61.6^{***}$ $p<0.001^{***}$ $p=0.092$ $p=0.220$
DTD $58.6^{***}$ $p=0.009^{**}$ $p=1.000$ $p=0.013^{*}$
EuroSAT $55.2^{***}$ $p=0.795$ $p=0.850$ $p=0.261$
Oxford Pets $63.5^{***}$ $p=0.417$ $p=1.000$ $p=0.055$